OpenStudy (helpmeee123):

Can someone help me with this plz? Plz show ur work .... A local high school collected $1590 from 321 people who attended a football game. The price of each adult admission is $6. People between the ages of 4-17 paid a children admission rate of $4. How many adult tickets and child tickets were sold that day?

OpenStudy (anonymous):

Name your unknown a= number of adult ticket c= number of kid ticket 6a + 4c = 1590 a + c = 321 Solve for a and c by using substitution a= 321-c now substitute 6(321-c) + 4 c = 1590 1926 - 6c + 4c = 1590 1950 - 1590 = 6c -4c 336 = 2c 336/2 = 2c/2 c= 168 a + c = 321 a+ 168 = 321 c= 321-168 = 153

OpenStudy (anonymous):

yea hes right